The length of the rectangular lot is 6 more then its widht. If the perimeter is 152 m. What are its dimensions?
Answer by addingup(3677)

2L+2W = Perimeter
and
L = W+6 substitute is the formula above:
2(W+6)+2W = 152
2W+12+2W = 152
4W = 140
W = 35
L = W+6 = 35+6 = 41
---------------------------------
Check:
2L+2W = 152
2(41)+2(35) = 152
82+70 = 152 Correct.
:
John